Question 11 Tries remaining: 2 Points out of 8.33 P Flag question Suppose that a price-searcher firm had consumers who were all identical to each other. The individual consumers demand function is givenby: 10 -5P. The firm decides to try a second-degree price discrimination scheme. The first 2 units will have a price of $1.60. After that, any units a consumer purchases will be only $1.20. The firm has a constant marginal cost of $1.00 per unit. Calculate the consumer surplus. (Do not include a $ sign in your response. Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

Given, Q = 10 - 5P

Inverse demand function, P = 2 - 0.2Q

We can plot the demand curve on the graph as shown below

4 CS 2

In the picture the consumer surplus is depicted we can calculate it as follows

Consumer Surplus = CS1 + CS2

Consumer Surplus = 0.5*(2-1.6)*2 + 0.5*(1.6-1.2)*(4-2)

Consumer Surplus = 0.4 + 0.4 = $ 0.80

Please contact if having any query thank you.